Circular No. 7037

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SCORPII 1998 W. Liller, Viña del Mar, Chile, writes: “A low-resolution CCD spec-

trogram obtained on Oct. 23.04 UT with a 0.20-m Schmidt camera (+ objective prism) shows a broad, intense Hα emission line at a level ∼ 7.6 times brighter than the surrounding continuum. This strong emission explains why the visually determined magnitudes are fainter (by roughly a magnitude) than those obtained with a CCD. On Oct. 23.003, I measured the magnitude with a CCD (+ a minus-infrared filter) to be 7.78.”

M. Fujii, Kurashiki, Okayama, Japan, reports that spectroscopic observations (0.28-m reflector; resolution ∼ 1 nm; instrumental resolution ∼ 450 km/s) on Oct. 24.42 UT show strong, broad Hα emission (FWHM ∼ 3800 km/s and strong, broad Hβ emission. The spectrum can be seen at

http://www1.harenet.ne.jp/~aikow/n sco1998.gif.

Visual magnitude estimates: Oct. 22.756 UT, 8.6 (B. Monard, Pretoria, South Africa); 22.91, 8.7 (J. G. de S. Aguiar, Campinas, S.P., Brazil); 23.41, 8.5 (A. Jones, Nelson, New Zealand); 23.912, 8.9 (Aguiar); 24.04, 9.0 (W. G. Dillon, Missouri City, TX).

SUPERNOVA 1998eg IN UGC 12133 M. Salvo, A. Pastorello, L. Rizzi, and V. Buondí, Asiago Astrophysical

Observatory, report that inspection of a preliminary spectrum (range 380– 700 nm) obtained on Oct. 20.45 UT with the 1.82-m telescope (+ AFOSC) shows that SN 1998eg (V ∼ 17.4) is a type-Ia supernova soon after maximum. The spectrum shows the 635.5-nm absorption of Si II typical of type-Ia supernovae. Given the redshift of the parent galaxy (7500 km/s), the expansion velocity deduced from the minimum of Si II (635.5 nm) is ∼ 10 500 km/s.

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1998eg was obtained by P. Berlind on Oct. 24.1 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), showing it to be a type-Ia supernova near maximum brightness. The age determined from a fit to the observed spectral features (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days after maximum light. The spectrum also exhibits redshifted narrow interstellar Na I D lines, implying significant extinction by dust in the host galaxy and yielding a recession velocity for the host galaxy of 7400 km/s. The expansion velocity of the supernova is 10 200 km/s, measured using the prominent Si II (rest 635.5 nm) feature.

1998 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green